• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:12:04 on February 7,2003

    Entry number 95173

    keyword=target oven RTD replaced, EPR fiber switched

    During the controlled access, Vince and I replaced oven RTD. We had to take out
    the middle cover as well as the top cover to get to the connector. The old one was broken and new one works fine so far.

    We also switched the longitudinal EPR fiber with the transverse EPR fiber. The
    longitudinal EPR was not working this morning. When Patricia tried EPR, the DC
    signal was very large. By switching the fiber we will get back to use two D2 filters
    which may help.